Show that the vectors u1 = (1, 1, 1), u2 = (1, 2, 3), u3 = (1, 5, 8) span R

3

.


1
Expert's answer
2021-10-22T01:16:22-0400

Solution:

To solve it show that an arbitrary vectorsl v=(a,b,c) "\in" R3 is a linear combination of "u_1,u_2" and "u_3" .

Let

"v=xu_1+yu_2+zu_3"

Hence,

"(a,b,c)=x(1,1,1)+y(1,2,3)+z(1,5,8)"

Resolve as

"x+y+z=a"

"x+2y+5z=b"

"x+3y+8z=c"

Form the equivalent matrix form of the equations

"\begin{bmatrix}
 1&1& 1&\vert&a \\
 1&2 & 5&\vert&b\\
1&3&8&\vert&c
\end{bmatrix}"

Reduced echelon form is

"\begin{bmatrix}
 1 & 1&1&\vert&a\\
 0 & 1&1&\vert&b-a\\
0&0&-1&\vert&a-2b+c
\end{bmatrix}"

From the reduced echelon form,

"-z=a-2b+c"

"z=-a+2b-c"

"y+4z=b-a"

"y=b-a-4z"

By substitution,

"y=b-a-4(-a+2b-c)"

"y=3a-7b+4c"

Also from the echelon form,

"x+y+z=a"

"x=a-y-z"

By substitution,

"x=a-(3a-7b+4c)-(-a+2b-c)"

"x=-a+5b-3c"

Therefore,

"(a,b,c)=(-a+5b-3c)(1,1,1)+(3a-7b+4c)(1,2,3)+(-a+2b-c)(1,5,8)"

Therefore,

"u_1,u_2,u_3" span "R^3"